RawlsK29:

Consider the following set of equations: Equation M: 3y = 3x + 6 Equation P: y = x + 2 Which of the following best describes the solution to the given set of equations? No solution One solution Two solutions Infinite solutions

RawlsK29:

@Bearclaws72 @Falconmaster

jhonyy9:

3y = 3x +6 y = x+2 like a first step try using the second equation ,so the y = x+2 value in the first equation and see what will get for x second step : using this x value in second equation will get easy the y value and so will result to you easy the right answer good luck